OpenStudy (anonymous):

Calc1 Optimization: A soup can is to be made in the shape of a cylinder, where the surface area of the label (the rectangle) is to be 100cm^2. Find the dimensions of the cylinder so that its volume is maximized, with the restriction that the radius r is 0

OpenStudy (anonymous):

v = pir^2h SA = 100cm^2 SA = r + h 100/(2pir) - r = h I think this is what you need I haven't done an optimization problem in a while though. Do you know where to go from here?

OpenStudy (anonymous):

sorry ignore the SA = r + h I was moving stuff around that should actually say SA = 2pir^2 + 2pirh

OpenStudy (anonymous):

I substitute in the H into the volume formula, then multiply by the denominator of H, then simplify. afterwards, I set it equal to 0 to find the optimal radius by isolating r (since according to the other problems I did, the only variable left should be r)

OpenStudy (anonymous):

Yes just think about it like this. When you are doing an optimization problem you are really solving for maxima and minima of a function. So you need to solve for the zeros of the function you are trying to find the optimal value for. Want to show me what you got?

OpenStudy (anonymous):

I'll get back to you on that in a bit

OpenStudy (anonymous):

Okay sounds good. Good luck.

OpenStudy (anonymous):

sorry I mean the zeros if the derivative of the function you are trying to find the optimal value for** basically what you said you were going to do

OpenStudy (anonymous):

of** not if haha it's late

OpenStudy (anonymous):

no problem, thanks for really helping me out today. I'm curious how you got: 100/(2pir) - r = h

OpenStudy (anonymous):

I factored 2pir out of the SA formula like this SA = 2pir(r+h) then I divided both side by 2pir then I subtracted r from both sides to get it in terms of H and I set the SA to the value you gave 100cm^2

OpenStudy (anonymous):

ok i see it was just algebra. I'll keep working

OpenStudy (anonymous):

Wait a second I just realized the SA is only for the rectangle not the full cylinder this changes things. It says the label only goes over the rectangle part.

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

The formula for SA I put is for the entire cylinder but your question only involves the rectangular part of the cylinder.

OpenStudy (anonymous):

what changes then?

OpenStudy (anonymous):

The formula for SA = 2pirh now so you have to sub h = 100/(2pir) instead I believe. Like I said I am rusty on this stuff tho.

OpenStudy (anonymous):

Here's what I got: SA = 2πrh 100 = 2πrh 100/(2πr) = h 50/πr = h Volume = hπr^2 = πr^2 * (50/πr) = r50 we hit a problem

OpenStudy (anonymous):

yeah sorry I messed up earlier one sec I think I can fix it

OpenStudy (anonymous):

(100-2pir^2)/(2pir) = h v =2pir^2((100-2pir^2)/(2pir)) v = (200pir^2- 2pir^4)/(2pir) never mind same issue as before I must be leading you in the wrong direction I am going to try and figure this out for a couple more minutes but I gotta go to sleep soon best of luck sorry I couldn't be more helpful.

OpenStudy (anonymous):

I'll work on it some more. thanks for trying

OpenStudy (anonymous):

Okay I think I got it after all I accidentally wrote v = 2pir^2h when I should have wrote v = pir^2h anyway v = ((100pir^2-2pi^2r^4)/(2pir)) v = (50r-pir^3) v' = 50 - 3pir^2 Now we find zeros

OpenStudy (anonymous):

0 = 50 - 3pir^2 r = (50/(3pi))^(1/2) sub that back into formula for volume without "h" and solve for volume to get (50((50/(3pi))^(1/2))-pi((50/(3pi))^(1/2))^3) V = 76.78 Now plug V back into formula this time with h and solve for h 76.78= pi(((50/(3pi))^(1/2))^2h h = 76.78/(pi(((50/(3pi))^(1/2))^2) h = 4.6 I think this is right, but as I am sure you could see I am pretty rusty with this. If you have the chance send me a message if it is right or wrong.